# Localization and entanglement characterization of edge states in HgTe quantum wells in a finite strip geometry

### v2

@@ IV.1 Analysis of near-gap state probability densities
[…]
@@
Figure 8: Probability density $|\Psi_{c,v}(k,y)|^{2}$ of the four 4-spinor near-gap states $\{\Psi_{1c},\Psi_{1v},\Psi_{2c},\Psi_{2v}\}$ (conduction in red and valence in blue) as a function of position $y\in[0,L],\,L=100$ nm, for several values of $k\in[0,\pi/a)$ (symmetric in $k$). The thickness of the probability density curves decreases with $k$, the thickest one corresponding to $k=0$. The maximum probability density at the edges $y=0,L$ corresponds to: $k\simeq\pm 0.21\,\mathrm{nm}^{-1}$ for valence band (blue) and $k=0$ for conduction band (red) staes. See later in Figure 12 for a 3D version.
@@
Figure 9: Component-wise probability densities of the four near-gap states $\{\Psi_{1c},\Psi_{1v},\Psi_{2c},\Psi_{2v}\}$, as a function of $y\in[0,L]$ ($L=100$nm), for three values of the momenta: $k_{x}=-0.005\,\mathrm{nm}^{-1}$ (dashed) $k_{x}=0.005\,\mathrm{nm}^{-1}$ (dotted) and $k_{x}=0\,\mathrm{nm}^{-1}$ (solid). Conduction in red and valence in blue colors. The spin polarization structure of the QSH edge states can be glimpsed.
@@
Figure 10: Spin-polarization of edge states in a strip geometry of width $L$: the spin $s$ (=1,=1$\uparrow$=1$\downarrow$=-1)↑ = 1 , ↓ = - 1 ) is aligned with the propagation direction given by ${k_{x}}$ so that the helicity $h=\mathrm{sign}(sk_{x})$ is $p=1$ for valence (blue) states at the edge $y=L$ and conduction (red) states at the edge $y=0$, and $p=-1$ for valence states at $y=0$ and conduction states at $y=L$.
Just looking at Hamiltonian spectra in Fig. 4 or density distributions in Fig. 8, it is difficult to establish a clear criterion for distinguishing between edge and bulk states as $k$ moves inside the interval $[0,\pi/a]$. In Ref. [6] a cutoff $k_{c}$, bounding the edge state region for the uncoupled tight-binding Hamiltonian, is proposed, above which edge states are not longer normalizable and touch the bulk (extended) states. A similar analysis for the SOC case becomes much more difficult and we have not found any additional result concerning this case in the literature. Here we propose a cutoff for the coupled case attending to a practical indicator of localization that has been utilized in a variety of settings, called the inverse participation ratio (IPR). This information measure is closely related to others like the Rényi entropy. Let us remind its definition.
@@ IV.2 IPR and localization in position space $y$ for a given momentum $k_{x}$
[…]
@@
Fig. 11 displays $\mathrm{IPR}_{\Psi}(k)$ for the first near-gap valence $\Psi_{1v}$ and conduction $\Psi_{1c}$ states (similar results for $\Psi_{2v}$ and $\Psi_{2c}$) as a function of $k_{x}$ for $L=100$ nm. Its shape is symmetric, as expected. Note that localization of valence and conduction states occurs here at different scales (roughly a factor of ten). $\Psi_{1v}$ shows maximum localization (maximum IPR) at ±$\pm k_{m}\simeq\pm 0.229\,\mathrm{nm}^{-1}$ and $\Psi_{1c}$ at ±$\pm k_{m}\simeq\pm 0.01\,\mathrm{nm}^{-1}$. Remember that a cutoff $k_{c}$ was defined in Fig. 1 for the analytic approach as the value of $k_{x}$ at which edge and bulk states merge, that is, when $k_{y}=0=-\mathrm{i}\lambda$. In the tight-binding approach, a reasonable definition of $k_{c}$ for a given near-gap state $\Psi$ is when $\text{IPR}_{\Psi}(k_{c})$ is comparable to $\text{IPR}_{\Psi}(0^{\pm})$, that is
@@
This definition requires in general to take the one-sided limit $k\to 0^{\pm}$ to avoid those cases in which there is a removable discontinuity of $\text{IPR}_{\Psi}(k)$ at $k=0$. The condition (25) gives $k_{c}\simeq 0.383\,\mathrm{nm}^{-1}$ for valence and $k_{c}\simeq 0.03\,\mathrm{nm}^{-1}$ for conduction first near-gap states as depicted in Fig. 11.
Figure 11: Localization $\text{IPR}_{\Psi}(k)$ of the first near-gap valence $\Psi_{1v}$ (upper panel, blue) and conduction $\Psi_{1v}$ (lower panel, red) in position space $y$ as a function of the momentum $k_{x}$ for a strip width of $L=100$ nm. Maximum $\text{IPR}_{1v}(k)$ is attained at ±$\pm k_{m}\simeq\pm 0.229\,\mathrm{nm}^{-1}$ and maximum $\text{IPR}_{1c}(k)$ is attained at ±$\pm k_{m}\simeq\pm 0.01\,\mathrm{nm}^{-1}$. Cutoff momentum $k_{c}$ is defined as $\text{IPR}_{\Psi}(k_{c})=\text{IPR}_{\Psi}(0^{\pm})$, which gives $k_{c}\simeq 0.383\,\mathrm{nm}^{-1}$ for valence and $k_{c}\simeq 0.03\,\mathrm{nm}^{-1}$ for conduction first near-gap states.
@@
Table 2: Edge state momentum cutoff $k_{c}$ for $\Psi_{1c}$ and $\Psi_{1v}$ as a function of the strip width $L=100$.
@@ IV.3 IPR and localization in momentum space $k_{x}$ for a given position $y$
[…]
Fig. 13 shows that near-gap states participate of less and less momenta $k$ as we approach the boundaries $y=0,L$ (higher IPR), since momentum is localized around ±$\pm k_{m}\simeq\pm 0.229\,\mathrm{nm}^{-1}$, as it was noted before in Fig 11. The second near-gap conduction $\Psi_{2c}$ and valence $\Psi_{2v}$ states also participate of less and less momenta $k$ (higher IPR) towards the center of the strip $y=L/2$.
@@ V.1 Reduced density matrix, spin probabilities and linear entropy
The $2\times 2$ reduced density matrix (RDM) $\varrho$ to the spin subsystem is obtained by taking the partial trace
@@
represent the probabilities of finding the electron with spin up or down, respectively, whereas the modulus of the off-diagonal elements
@@
In Fig. 14 we plot the probabilities $P_{\Psi}(\uparrow)$ and $P_{\Psi}(\downarrow)$ for the first conduction $\Psi_{1c}$ and valence $\Psi_{1v}$ near-gap states as a function of $(k,y)$ in the entire Brillouin zone $[-\pi/a,\pi/a]$. Lighter colors represent higher probability zones. Probability densities are unbalanced at the boundaries $y=0,L$ for low momenta $k$, depending on the propagation direction given by the sign of $k$. This is again a reflection of the existence of counterpropagating modes of opposite spin at the edges (helical edge modes), which is consistent with the experimental confirmation in [22] that the transport in the edge channels is spin polarized. This is so because, at each edge of the strip, the sign $\text{sign}(ks)$ is well defined, being $+1$ for the $y=0$ edge and $-1$ for the $y=L$ edge.
@@ V.2 Schlienz & Mahler entanglement measure
VI Conclusions
We have used QI theory concepts like IPR, RDM and entanglement entropies, as an interesting “microscope” to reveal details of the internal structure of HgTe QW near-gap states with SOC (induced by the bulk and structural inversion asymmetries) in a finite strip geometry of width $L$, proposing a criterion for a momentum cutoff $k_{c}$ differentiating between edge- and bulk-states region behavior attending to IPR localization measures. To do this, we have considered a tight-binding Hamiltonian describing the low energy effective theory. Quantitative information on the near-gap energy spectrum and wave-functions is extracted from a numerical Hamiltonian diagonalization approach, which is complemented by a previous analytic study of the uncoupled (without SOC) case.

### v3

@@ IV.1 Analysis of near-gap state probability densities
and normalize them according to $\int_{0}^{L}dy|\Psi(k,y)|^{2}=1$ (in the continuum limit $a\to 0$ of small lattice constant).
@@
Figure 8: Probability density $|\Psi_{c,v}(k,y)|^{2}$ of the four 4-spinor near-gap states $\{\Psi_{1c},\Psi_{1v},\Psi_{2c},\Psi_{2v}\}$ (conduction in red and valence in blue) as a function of position $y\in[0,L],\,L=100$ nm, for several values of $k\in[0,\pi/a)$ (symmetric in $k$). The thickness of the probability density curves decreases with $k$, the thickest one corresponding to $k=0$. The maximum probability density at the edges $y=0,L$ corresponds to: $k\simeq\pm 0.21\,\mathrm{nm}^{-1}$ for valence band (blue) and $k=0$ for conduction band (red) states. See later in Figure 12 for a 3D view.
@@
Figure 9: Component-wise probability densities of the four near-gap states $\{\Psi_{1c},\Psi_{1v},\Psi_{2c},\Psi_{2v}\}$, as a function of $y\in[0,L]$ (L=100L=100nm), for three values of the momenta: $k_{x}=-0.005\,\mathrm{nm}^{-1}$ (dashed) $k_{x}=0.005\,\mathrm{nm}^{-1}$ (dotted) and $k_{x}=0\,\mathrm{nm}^{-1}$ (solid). Conduction in red and valence in blue colors. The spin polarization structure of the QSH edge states can be glimpsed.
@@
Figure 10: Spin-polarization of edge states in a strip geometry of width $L$: the spin $s$ ($\uparrow=1,\downarrow=-1)$ is aligned with the propagation direction given by $k_{x}$} so that the helicity $h=\mathrm{sign}(sk_{x})$ is $p=1$ for valence (blue) states at the edge $y=L$ and conduction (red) states at the edge $y=0$, and p=1p=-1 for valence states at $y=0$ and conduction states at $y=L$.
Just looking at Hamiltonian spectra in Fig. 4 or density distributions in Fig. 8, it is difficult to establish a clear criterion for distinguishing between edge and bulk states as $k$ moves inside the interval $0,\pi/a]$. In Ref. [6] a cutoff $k_{c}$, bounding the edge state region for the uncoupled tight-binding Hamiltonian, is proposed, above which edge states are not longer normalizable and touch the bulk (extended) states. A similar analysis for the SOC case becomes much more difficult and we have not found any additional result concerning this case in the literature. Here we propose a cutoff for the coupled case attending to a practical indicator of localization that has been utilized in a variety of settings, called the inverse participation ratio (IPR). This information measure is closely related to others like the Rényi entropy. Let us remind its definition.
@@ IV.2 IPR and localization in position space $y$ for a given momentum $k_{x}$
In general, the IPR measures the spreading of the expansion of a normalized vector $|\Psi\rangle=\sum_{n=1}^{N}c_{n}|n\rangle$ in a given basis $\{|n\rangle,n=1,\dots,N\}$. For a normalized vector $|\Psi\rangle$, the probabilities $p_{n}=|c_{n}|^{2}$ to find $|\Psi\rangle$ in a given basis state $|n\rangle$ fulfil $\sum_{n=1}^{N}p_{n}=1$. The inverse participation ratio of $\Psi$ is defined as
$\mathrm{IPR}_{\Psi}=\sum_{n=1}^{N}p_{n}^{2}=\sum_{n=1}^{N}|c_{n}|^{4}.$
It reduces to the square of the $\alpha$-norm $\|\Psi\|_{\alpha}=(\sum_{n=1}^{N}p_{n}^{\alpha})^{1/\alpha}$ for $\alpha=2$. Moreover, it is also related to the Rényi entropy of order $\alpha$
$R^{\alpha}_{\Psi}=\frac{1}{1-\alpha}\log\sum_{n=1}^{N}p_{n}^{\alpha},$
for $\alpha=2$. Rényi entropy generalizes the quantum mechanical $\alpha\to 1$ von Neumann entropy. The smallest value $\mathrm{IPR}_{\psi}=1/N$ corresponds to an equally weighted superposition $|c_{n}|=1/\sqrt{N},\forall n$ (fully delocalized state), whereas the highest value $\mathrm{IPR}_{\psi}=1$ corresponds to a perfectly localized state, that is, $c_{n}=\delta_{n,n_{0}}$ for some $n_{0}$.
The concept can be similarly extended to the representation of the wavevectors in position, momentum or other spaces, so as to study the spreading, or the localization, of the vector in the given space basis. For example, for the case of a free particle in a box $y\in[0,L]$, the wave function $\psi_{m}(y)=\sqrt{2/L}\sin(m\pi y/L)$, normalized according to $\int_{0}^{L}dy|\psi_{m}(y)|^{2}=1$, has an $\mathrm{IPR}_{m}=\int_{0}^{L}dy|\psi_{m}(y)|^{4}=3/(2L)$. For example, for a strip width of $L=100$ nm, we have $\mathrm{IPR}_{m}=0.015$.
In our case, the IPR of 4-spinor $\Psi(k,y)$, normalized as $\int_{0}^{L}dy|\Psi(k,y)|^{2}=1$, in position $y$ space for each value of the momentum $k$ is given by
$\mathrm{IPR}_{\Psi}(k)=\int_{0}^{L}dy|\Psi(k,y)|^{4},$ (23)
@@
Fig. 11 displays $\mathrm{IPR}_{\Psi}(k)$ for the first near-gap valence $\Psi_{1v}$ and conduction $\Psi_{1c}$ states (similar results for $\Psi_{2v}$ and $\Psi_{2c}$) as a function of $k_{x}$ for $L=100$ nm. Its shape is symmetric, as expected. Note that localization of valence and conduction states occurs here at different scales (roughly a factor of ten). $\Psi_{1v}$ shows maximum localization (maximum IPR) at $\pm k_{m}\simeq\pm 0.229\,\mathrm{nm}^{-1}$ and $\Psi_{1c}$ at $\pm k_{m}\simeq\pm 0.01\,\mathrm{nm}^{-1}$.
Remember that a cutoff $k_{c}$ was defined in Fig. 1 and Eq. (14) for the analytic approach as the value of $k_{x}$ at which edge and bulk states merge. That is, after the replacement $k_{y}\to-\mathrm{i}\lambda$ done in the Hamiltonian $H(k_{x},k_{y})$ at the beginning of Section III.1 to look for edge localized solutions $\Psi(y)\sim e^{\lambda y}$ of the Schrödinger equation, the criterion for edge-bulk transmutation was $\lambda=0=\mathrm{i}k_{y}$, providing the cut-off values of $k_{x}$ given in equation (14).
In the tight-binding approach, our proposal for a proper definition of $k_{c}$ for a given near-gap state $\Psi$ is when $\text{IPR}_{\Psi}(k_{c})$ is comparable to $\text{IPR}_{\Psi}(0^{\pm})$, that is
@@
This definition requires in general to take the one-sided limit $k\to 0^{\pm}$ to avoid those cases in which there is a removable discontinuity of $\text{IPR}_{\Psi}(k)$ at $k=0$. Our proposal for $k_{c}$ is physically justified by the fact that low momentum $k\to 0$ states correspond to translationally invariant (highly delocalized, low IPR) configurations. Therefore, for near-gap (low momentum) states $\Psi$, a reasonable definition of $k_{c}$ is when $\text{IPR}_{\Psi}(k_{c})$ attains low values comparable to $\text{IPR}_{\Psi}(0^{\pm})$.
The condition (25) gives $k_{c}\simeq 0.383\,\mathrm{nm}^{-1}$ for valence and $k_{c}\simeq 0.03\,\mathrm{nm}^{-1}$ for conduction first near-gap states for $L=100$ nm, as depicted in Fig. 11.
Figure 11: Localization $\text{IPR}_{\Psi}(k)$ of the first near-gap valence $\Psi_{1v}$ (upper panel, blue) and conduction $\Psi_{1v}$ (lower panel, red) in position space $y$ as a function of the momentum $k_{x}$ for a strip width of $L=100$ nm. Maximum $\text{IPR}_{1v}(k)$ is attained at $\pm k_{m}\simeq\pm 0.229\,\mathrm{nm}^{-1}$ and maximum $\text{IPR}_{1c}(k)$ is attained at $\pm k_{m}\simeq\pm 0.01\,\mathrm{nm}^{-1}$. Cutoff momentum $k_{c}$ is defined as $\text{IPR}_{\Psi}(k_{c})=\text{IPR}_{\Psi}(0^{\pm})$, which gives $k_{c}\simeq 0.383\,\mathrm{nm}^{-1}$ for valence and $k_{c}\simeq 0.03\,\mathrm{nm}^{-1}$ for conduction first near-gap states.
@@
Note that these cutoff $k_{c}$ values differ from those obtained by formula (14) which, for example, give $k_{c}\simeq 0.004\,\mathrm{nm}^{-1}$ for conduction first near-gap states for $L=100$ nm. This is because, in addition to the fact that we are using a different (numeric, tight-binding) approach, we are also considering full SOC (BIA plus SIA) interaction, which was absent in the analytic approach of Section III.1.
@@
Table 2: Edge state momentum cutoff $k_{c}$ for $\Psi_{1c}$ and $\Psi_{1v}$ as a function of the strip width for $L=100,200$ and $250$ nm.
@@
The probability densities $|\Psi_{1c,1v}(k_{c},y)|^{2}$, in magenta color in Fig. 12, correspond to the flattest ones in the real $y$-space, that is, the quantity
$\textrm{Max}_{y\in[0,L]}|\Psi(k,y)|^{2}-\textrm{Min}_{y\in[0,L]}|\Psi(k,y)|^{2}$ (26)
exhibits a minimum value for $k\simeq k_{c}$. Note that this criterium approximately coincides with that of Eq. (25).
@@ IV.3 IPR and localization in momentum space $k_{x}$ for a given position $y$
For completeness, we also analyze the spreading of the expansion of near-gap states in momentum space $k$ for a given position $y$. To do that, we have to normalize 4-spinors as $\int_{-\pi/a}^{\pi/a}dk|\Psi(k,y)|^{2}=1$ and define
$\mathrm{IPR}_{\Psi}(y)=\int_{-\pi/a}^{\pi/a}dk|\Psi(k,y)|^{4}.$ (27)
Fig. 13 shows that near-gap states participate of less and less momenta $k$ as we approach the boundaries $y=0,L$ (higher IPR), since momentum is localized around $\pm k_{m}\simeq\pm 0.229\,\mathrm{nm}^{-1}$, as it was noted before in Fig 11. The second near-gap conduction $\Psi_{2c}$ and valence $\Psi_{2v}$ states also participate of less and less momenta $k$ (higher IPR) towards the center of the strip $y=L/2$.
@@ V.1 Reduced density matrix, spin probabilities and linear entropy
The $2\times 2$ reduced density matrix (RDM) valued $\varrho(k,y)$ to the spin subsystem is obtained by tracing out the orbital $E,H$ degrees of freedom
@@
represent the probabilities of finding the electron with spin up or down at each point $(k,y)$, respectively, whereas the modulus of the off-diagonal elements
@@
In Fig. 14 we plot the probabilities $P_{\Psi}(\uparrow)$ and $P_{\Psi}(\downarrow)$ for the first conduction $\Psi_{1c}$ and valence $\Psi_{1v}$ near-gap states as a function of $(k,y)$ in the entire Brillouin zone $\pi/a,\pi/a]$. Lighter colors represent higher probability zones. Probability densities are unbalanced at the boundaries $y=0,L$ for low momenta $k$, depending on the propagation direction given by the sign of $k$. This is again a reflection of the existence of counterpropagating modes of opposite spin at the edges (helical edge modes), which is consistent with the experimental confirmation in [27] that the transport in the edge channels is spin polarized. This is so because, at each edge of the strip, the sign $\text{sign}(ks)$ is well defined, being +1+1 for the $y=0$ edge and 1-1 for the $y=L$ edge.
@@ V.2 Schlienz & Mahler entanglement measure
The conclusion here is that the behavior of quantum correlations in HgTe QWs appears to be independent of the specific entanglement measure employed.
VI Correlation functions for real-space partitions
In this Section we propose a real-space partition $y\in[0,L_{1}]\cup[L_{1},L]$ (that is, a bipartition of the system along the $y$-direction) to probe edge versus bulk behavior. Other real-space partitions appear in the literature [12, 13]. Let $|\Psi\rangle$ a $4N$ component spinor state; for example, an eigenstate of the tight-binding Hamiltonian (18) for a strip width $L=aN$ and momentum $k=k_{x}$. The two-point correlation function for this state is
$C_{\Psi}(k,m,n)=\langle\Psi|c^{\dagger}_{k,m}c_{k,n}|\Psi\rangle,$ (39)
where $c_{k,n}$ and $c^{\dagger}_{k,n}$ are the annihilation and creation operators with momentum $k$ at lattice site $n$ which define the tight-binding Hamiltonian (18). For the real-space partition $y\in[0,L_{1}]\cup[L_{1},L]$, the space correlation function is then
$\overline{C}_{\Psi}(k)=\frac{\sum_{m=0}^{N_{1}}\sum_{n=N_{1}+1}^{N}C_{\Psi}(k,m,n)}{\sqrt{(N_{1}(N-N_{1})}},$ (40)
with $N_{1}=L_{1}/a$. In the continuum limit (small lattice constant) it could be written as
$\overline{C}_{\Psi}(k)=\frac{\int_{0}^{L_{1}}\int_{L_{1}}^{L}dydy^{\prime}\bar{\Psi}(k,y)\Psi(k,y^{\prime})}{\sqrt{(L_{1}(L-L_{1})}},$ (41)
for a normalized probability density (22). Figures 19 and 20 show plots of the correlation function (modulus and argument)
$\overline{C}_{\Psi}(k)=|\overline{C}_{\Psi}(k)|e^{\mathrm{i}\Theta_{\Psi}(k)}$ (42)
for the space-partition $L_{1}=L/2$ and the first near-gap valence $\Psi_{1v}$ (blue) and conduction $\Psi_{1c}$ (red) states, as a function of the momentum $k=k_{x}$ for a strip width of $L=100$ nm. Comparing with Fig. 11, we see that higher IPR values correspond to lower space-correlation modulus $|\overline{C}_{\Psi}(k)|$ values. Moreover, space-correlation arguments/angles $\Theta_{\Psi}(k)$ attain their maximum absolute values in the edge-state (high IPR) region, exhibiting positive (resp. negative) values for negative (resp. positive) momenta $k$. To better compare both information measures, we overprint the maximum IPR values at $\pm k_{m}\simeq\pm 0.229\,\mathrm{nm}^{-1}$ for valence (blue) and at $\pm k_{m}\simeq\pm 0.01\,\mathrm{nm}^{-1}$ for conduction (red) states, respectively. We also mark the momentum cutt-offs $k_{c}\simeq 0.383\,\mathrm{nm}^{-1}$ for valence (blue) and $k_{c}\simeq 0.03\,\mathrm{nm}^{-1}$ for conduction (red) first near-gap states obtained from the IPR criterion (25). In general, we could say that edge states participate of fewer momenta $k$ (higher IPR) than bulk states; they also have lower correlation modulus (but higher correlation arguments in absolute value) than bulk states for the real-space partition discussed here.
Figure 19: Correlation fucntion $C_{\Psi}(k)$ (modulus $|C_{\Psi}(k)|$ and argument $\Psi_{\Psi}(k)$) of the first near-gap valence state $\Psi_{1v}$ for a real-space partition $y\in[0,L/2]\cup[L/2,L]$ as a function of the momentum $k$. Maximum $k_{m}$ and cut-off $k_{c}$ momenta of Fig. 11 (upper panel) are marked with vertical grid lines to match IPR with correlation functions.
Figure 20: Correlation fucntion $C_{\Psi}(k)$ (modulus $|C_{\Psi}(k)|$ and argument $\Psi_{\Psi}(k)$) of the first near-gap conduction state $\Psi_{1c}$ for a real-space partition $y\in[0,L/2]\cup[L/2,L]$ as a function of the momentum $k$. Maximum $k_{m}$ and cut-off $k_{c}$ momenta of Fig. 11 (lower panel) are marked with vertical grid lines to match IPR with correlation functions.
VII Conclusions and outlook
We have used QI theory concepts like IPR, RDM, entanglement entropies and space correlations, as an interesting “microscope” to reveal details of the internal structure of HgTe QW near-gap states with SOC (induced by the bulk and structural inversion asymmetries) in a finite strip geometry of width $L$, proposing a criterion for a momentum cutoff $k_{c}$ differentiating between edge- and bulk-states region behavior attending to IPR localization measures. To do this, we have considered a tight-binding Hamiltonian describing the low energy effective theory. Quantitative information on the near-gap energy spectrum and wave-functions is extracted from a numerical Hamiltonian diagonalization approach, which is complemented by a previous analytic study of the uncoupled (without SOC) case.